OpenStudy (anonymous):

I need to solve this equation 2x over 5 + 5x=4 over 3

OpenStudy (anonymous):

\[2x/5+5x=4/3\] \\[[x(2/5+5)=4/3\]\]

OpenStudy (anonymous):

\[x(27/5)=4/3\]

OpenStudy (anonymous):

\[x=20/(27*3\]

OpenStudy (anonymous):

\[x=20/81\]

OpenStudy (anonymous):

Where did you get the x=20? by 4*5?

OpenStudy (anonymous):

you can simplify this equation by multiplying each side by 3: 6x/(5+5x)=4 factor out a 5 in the denominator: 6x/(5(1+x))=4 multiply both sides by 5: 6x/(1+x)=20 multiply by (1+x): 6x=20+20x subtract 20x: 6x -20x=20 -14x=20 divide by -14 x=20/-14=-10/7

OpenStudy (anonymous):

20/81 is not correct

OpenStudy (anonymous):

i failed !

OpenStudy (anonymous):

I got it...

OpenStudy (anonymous):

The book has 20/81 for an answer

OpenStudy (anonymous):

hi5

OpenStudy (anonymous):

fan me !

OpenStudy (anonymous):

ill write it out more clearly wait

OpenStudy (anonymous):

i must have written the equation wrong i thought it was 2x/(5+5x)=(4/3)

OpenStudy (anonymous):

Thank you soo much!!!

OpenStudy (anonymous):

@stair.. Yes, I see the 6x instead of 2x.

OpenStudy (anonymous):

\[2x/5+5x=4/3\] \[x(2/5+5)=4/3\] \[x(27/5)=4/3\] \[x=4/3*5/27\] \[x=20/81\]

OpenStudy (anonymous):

Thanks again!!

OpenStudy (anonymous):

no problem.